OpenStudy (anonymous):

A statue is made up of iron and nickel in a ratio of 2:7 . If the statue weighs 81 pounds, what is the weight of the nickel in the statue ?

OpenStudy (anonymous):

@UsukiDoll can you help me out ?

OpenStudy (ccswims):

2+7 = 9 81/9 = 9

OpenStudy (ccswims):

multiply 2 by 9 and 7 by 9

OpenStudy (ccswims):

then you get the weight of the nickel

OpenStudy (usukidoll):

sorry I was helping with another question.. I'll do verification in a bit.

OpenStudy (anonymous):

@cc swims I did it

OpenStudy (anonymous):

I got 18 & 63

OpenStudy (mathstudent55):

If the statue has a ratio of iron to nickel of 2:7, it could have: iron nickel 2 * 1 lb + 7 * 1lb 2 * 2 lb + 7 * 2 lb 2 * 3 lb + 7 * 3 lb 2 * 4 lb + 7 * 4 lb All the above are in a 2:7 ration, but we don't know which number to multiply both parts of the ratio by, so we call that unknown x. The ratio is 2x : 7x Since the total weight is 81 lb, we write the equation 2x + 7x = 81 9x = 81 x = 9 Then 2x = 2(9) = 18, and 7x = 7(9) = 63 The statue has 18 lb of iron and 63 lb of nickel.

OpenStudy (anonymous):

Thanks @mathstudent55 :)

OpenStudy (mathstudent55):

You're welcome.